What are the deltas of a call option and a put option with the following characteristics? (A negative answer should be indicated by a minus sign. Do not round intermediate calculations and round your answers to 4 decimal places, e.g., 32.1616.)

Stock price = $49
Exercise price = $45
Risk-free rate = 3.2% per year, compounded continuously
Maturity = 8 months
Standard deviation = 54% per year

Call option delta __________

Put option delta __________
Answer #1

Given, Stock Poice = $49 Exercise Price = $45 Risk-free rate= 3.2% per year, compounded continuously Maturity = 8 months StanS = Exercise Price Strike Price ka y Os Standard Dewlation t= time to maturity t SERTA So, according to formula; (da)= min )=(-0.0852) + ( 0.1185) 0.54 (0.81649) z 0.0333 0.4409 0.075 So, 101) - 0.075 . Delta= N(d 1) = 0.5299. as, value of N(0.07)=0- Delta for call option = 0.5299 Now, Delta for put option = N(d 1) - 1 20.5299- 1 ; -0,470) Answer